OpenStudy (anonymous):

can anyone write this in a formula? (x^4)^1/2

OpenStudy (mathstudent55):

\( (a^m)^n = a^{mn} \)

OpenStudy (anonymous):

can you type the one yo typed before? @mathstudent55

OpenStudy (mathstudent55):

\( a^{\frac{1}{n}} = \sqrt[n]{a} \)

OpenStudy (anonymous):

thank you

OpenStudy (mathstudent55):

wlcm